# Hand1000: Generating Realistic Hands from Text with Only 1,000 Images

### v2

@@ Conclusion
We have presented a novel method based on the diffusion model to generate realistic hand images from text using only 1,000 images. Through quantitative and qualitative experiments, we demonstrate the superiority and effectiveness of our approach. In particular, we show that it is significant to incorporate hand gesture information and optimize text embeddings for accurate hand image generation. As text-to-image generation continues to evolve, the methods and insights presented in this paper pave the way for more accurate and reliable visual representations, particularly in the challenging domain of human hands. In the future, we plan to build upon these findings to refine and expand the capabilities of text-to-image models across a broader range of applications.

### v3

@@ Conclusion
We have presented a novel method based on the diffusion model to generate realistic hand images from text using only 1,000 images. Through quantitative and qualitative experiments, we demonstrate the superiority and effectiveness of our approach. In particular, we show that it is significant to incorporate hand gesture information and optimize text embeddings for accurate hand image generation. As text-to-image generation continues to evolve, the methods and insights presented in this paper pave the way for more accurate and reliable visual representations, particularly in the challenging domain of human hands. While encouraging, this work is limited by the lack of exploration into simultaneous target gesture generation and hand-object interaction, primarily due to the unavailability of datasets, which will be our future work.
We would also like to gratefully acknowledge Bitdeer.AI for providing the cloud services and computing resources that were essential to this work.